From: Mythes, Croyances et Religions, vol. 18, 2000, p. 89-97.

A Childhood, The Gospel Singer, and Writing: Harry Crews’s Rituals of Exorcism

Anne FOATA Université Marc-Bloch, Strasbourg

“It has always seemed to me that I was not so much born into this life as I awakened to it. I remember very distinctly the awakening and the morning it happened. It was my first glimpse of myself,” Harry Crews writes in his autobiography A Childhood (47).

Crews was not quite five on that momentous morning when he awoke to his self. Within one year, however, his particular “postage stamp” of Southern Georgia, the family and kin he came from, the events that befell him and the mishaps he had to endure, were to mark, indeed to scar, him forever. They were to shape the man he has become and the kind of fiction he has written so far.

Nothing, not even the success and recognition that has come to the writer and the teacher of fiction at the University of Florida over the past thirty years, has been able to shake off the terrible burden laid upon the five-year-old boy by his uncongenial environment. Nothing, it seems, has allowed him to escape from himself and to be set free from the curse of violence and disaster branded on his flesh and spirit.

HARRY CREWS’S RITUALS OF EXORCISM

To read Crews’s 1978 “biography of a place” (the subtitle of A Childhood) from the vantage point of the turning century is thus to be handed the key to his entire work so far, fiction and nonfiction alike; and this holds not only for the first batch of novels that marked his entrance in the world of letters and culminated with the actual “biography,” but also for the second group that started appearing in the late eighties after Crews’s interlude of writing for the magazines and for Hollywood, when he was already well established as a novelist.

What Harry Crews was to retain from his devastated childhood in rural Georgia is best expressed by a character in his first novel, The Gospel Singer, who ruefully remarks that “there is nothing so predictable as the ritual of catastrophe and tragedy” (Gospel 198), a statement in which both terms are of equal importance, “the ritual” no less portentous than “the catastrophe and tragedy.”

Crews’s kith, kin and neighbors in Bacon County were the impoverished tenant farmers of the Depression years, riddled with the diseases and deformities that come with malnutrition and lack of medical care, crippled by the various accidents that are the poor farmer’s lot. They were also uneducated and subject to all kinds of superstitions, their women longsuffering and abused by the hard work on the farms, the large families, and their husbands’ and fathers’ heavy drinking. Asked in a 1972 interview if he would rather identify himself with the South of Caldwell than with the South of the Agrarians, of Faulkner or Welty, Crews assented, but he reminded his interviewer that Caldwell’s father was a minister and his family had enough to eat (Getting Naked 29); he added that things had not changed that much: his world was still, in 1972,“the southern landscape of Georgia, full of hookworms and poverty” (ibid. 32).

Crews’s was a world where one had to survive by any means and at any cost, where, for instance, the friend of a deceased man would steal all the meat stored in the widow’s smokehouse the night after the burial. Crews, who lost his father in 1937, when he was not yet two years old, writes that his mother knew who the thief was and that he was not only a friend but a close friend. But “[i]t was a hard time in that land, and a lot of men did things for which they were ashamed and suffered for the rest of their lives. [T]hey did them because of hunger and sickness and because they could not bear the sorry spectacle of their children dying from lack of a doctor and their wives growing old before they were thirty” (Childhood 42–43).

ANNE FOATA

It was also a world, Crews writes, that left people no margin for error or bad luck, where disasters compounded each other in a fatal escalation. “[W]hen something went wrong, it amost always brought something else down with it” (ibid. 40). This Crews illustrates by a mishap that occurred to him when he was a mere toddler and his mother was scrubbing the floor with homemade lye. Spotting the two yearling cows that they had ever owned about to drink from the barrel of lead poisoning with which her husband was spraying the tobacco nearby, she abandoned brush and lye and ran toward the field. Halfway there she heard Harry’s scream and knew what had happened. When the family returned from the doctor’s office in the nearby town (they had no car then, only a wagon and a mule), they also knew what was awaiting them: the spectacle of the two yearling cows lying dead and “already stiff” by the lead poisoning barrel (ibid. 39). Thinking about the episode as it was later recounted to him by his mother (as he had not yet awakened to his self), Crews muses how typical it was, and how tragic, of the world he grew up in, “a world in which survival depended on raw courage, a courage born out of desperation and sustained by a lack of alternatives” (ibid. 40).

Nor was the young Crews spared this kind of mishap as he grew up, and here we come again to that momentous fifth year of his age which marked the beginning of his conscious life and initiated the pattern of adversity and estrangement that was to be Crews’s crucial experience in life.

On the night of August 7, 1940, which, Crews writes, was exactly three months after his fifth birthday, he was taken by a high fever while his legs slowly bent at the knees with his heels getting closer and closer “to the cheeks of [his] buttocks” (ibid. 77). The pain was excruciating. And so he remained a couple of weeks in deep anguish and terror, with no one even knowledgeable enough to put a name on his affliction or to ease his pain: the two country doctors who quickly abandoned the case, a faith healer who spouted verses from Ezekiel, passing gypsies who sold his father ten dollars’ worth of herbs and stole a brood sow, and a tongue-speaking uncle who fell on him in a trance. The little boy’s only comfort as it were came from Auntie, the old black woman who lived on the Crews farm and had been born a slave, who managed to persuade him that “in this world there was much more to worry about than merely being crippled” (ibid. 83), a bitter truth that laid the groundwork of the future writer’s extensive and highly emblematic use of freaks.

Stared at by an uninterrupted string of gawking relatives and strangers who came from as far as the neighboring counties, Crews writes: “I felt how lonely and savage it was to be a freak” (ibid. 79), and a few pages further: “Right there, as a child, I got to the bottom of what it means to be lost, what it means to be rejected by everybody… and everything you ever thought would save you” (ibid. 83), knowing too by some kind of inbred knowledge that this was the condition of the world, the sheer mystery and terror of living.

Crews eventually recovered, but he was still not six years old when he had to endure yet another catastrophe which strengthened the sense of alienation he already so keenly felt. This is recounted in chapter 9 of A Childhood. It was at hog-slaughtering time the following winter, with all the neighbors assembled in the Crewses’ farm to lend a hand. The young Harry was playing “pop-the-whip” with the other children, a game in which everyone holds hands in a line and runs fast until the leader turns sharply, literally popping loose the last child on the line and sending him flying from his playmates. And, of course, being the last child on the line, the young Crews (who should have known better than play this crazy game on his newly-recovered legs), popped loose and went sailing into the steaming vat of boiling water in which the hogs were scalded.

Somebody fished him out and set him on his feet. He did not fall down, but looking at the horrified faces around him, he “saw” in them that he was dead.

“In memory I stand there alone with the knowledge of death upon me. [...] I reached over and touched my right hand with my left, and the whole thing came off like a wet glove. I mean, the skin of the top of the wrist and the back of my hand, along with the fingernails, all just turned loose and slid on down to the ground. I could see my fingernails lying in the little puddle my flesh made on the ground in front of me” (ibid. 113).

The skin of his back came off with his shirt, what was left on his “cooked” body with his overalls. He still had not fallen; he stood there “participating in [his] own butchering”. His pain meanwhile had “turned into something words cannot touch” (ibid.). The coup de grâce was dealt to him when out of ignorance and panic, they wrapped him in a blanket to drive him the sixteen miles to the nearby town of Alma.

This, Crews writes, was his “second major hurt back to back” in the same year, and it was not to be his last, as his years with the Marines and his wild motorcycling tour around the continent would provide their own share, but our concern here is with Crews’s annus horribilis when, aged five, according to his own admission, he came to the full consciousness of his being and of the world around him, of its unfathomable mystery and terror.

To assuage the terror, to find some defense against it and to make some sense of his life, the young Crews, still in the fifth year of his age, started to make up stories, to “fabricate,” as he says, “to spin a web of fantasy” (ibid. 54) around the world he experienced every day. Choosing random characters from the pages of the Sears, Roebuck catalogue, Harry and his black playmate, the son of the tenant farmers who lived on the farm, set to creating stories about them, inventing relationships between them, mostly violent ones, of those that included family feuds and ended in blood. For, at five, and already well in the know of all the couplings around him, animal and human, Crews also knew that if the smooth, well-dressed young man they spotted on one page of the catalogue was “fooling” with the pretty girl they found on another, “the man in his middle years, dressed in a hunting jacket and wading boots” (ibid. 55), whom they spotted on still another page and made out to be the girl’s father, would stop all that “messing around” with the guns and knives that were on display in a further section of the catalogue. They sometimes embroidered on the story by having the whole male kin lending a hand to the father and laying traps for the young man, and even refined it by having the girl’s complicity in the whole bloody affair. “Before it was over,” Crews writes, “the entire Wish Book (as the catalogue was called in the country) was filled with feuds of every kind and violence, maimings, and all the other vicious happenings of the world” (ibid. 57).

Another noteworthy thing about the catalogue, in the light of all the deformed characters that were to take up the space of his fiction, is that it showed people who were not only smiling and beautiful but whole, that is, unscarred, unmaimed, in a word, perfect. And as young as he was, Crews knew it was a lie, “that under those fancy clothes there had to be scars ... swellings and boils of one kind or another because there was no other way to live in the world” (ibid. 54).

Of those “vicious happenings of the world” that the boy devised for his first fantasies, the Crews family, meanwhile, was getting more than their normal share, and an increasingly uncontrollable one, on account of the stepfather’s drinking and the ensuing violence. It was during the winter of Harry’s fifth birthday that the sitation finally got out of hand. Life was becoming a nightmare, and the young Crews was well acquainted with nightmares and their unspeakable terror. “[I]t occurred to me for the first time that being alive was like being awake in a nightmare. I remember saying aloud to myself: ’Scary as a nightmare. Jest like being awake in a nightmare’”(ibid. 108).

Again, in the middle of the tension and terror of his daily life, of the mishaps of all kinds that were his daily bread, the young Crews was well aware that these were nothing compared to the torments of eternal damnation that God would eventually wreak upon them. He was familiar with them; they were preached to him during every Sunday service. And so it came to pass that to ease the tension of his mother and brother while they were waiting for their father, Crews, one night when he was five, improvised his own sermon and preached “about hell and God and heaven and damnation and the sorry state of the human condition” (ibid. 65). The adult Crews comments:

In all the churches, you smelled the brimstone and the sulfur and you felt the fire and you were made to know that because of what you had done in your life, you were doomed for ever. Unless somehow, somewhere, you were touched by the action of mercy and the Grace of God. But you could not, you must not, count on the Grace of God. It probably would not come to you because you were too sorry (ibid. 65-66; emphasis added).

The certainty embedded in those last two sentences may well encapsulate Crews’s whole “philosophy” of life (although Crews would certainly object to the word); they also foretell most of the contents of his novels and nonfictional stories; they explain most of his characters, their innate sorriness, their ever renewed hope in love, salvation… God, and their no less ever renewed disillusionment. Crews, at five, together with the gift of self-awareness, was given this poisoned glimpse into the human condition.

All of Crews’s characters do indeed hanker after something that might or might not be the Grace of God, but they also know that they are too sorry to ever be worthy of it. They inhabit a doomed world replete with all kinds of wretchedness that they are powerless to understand and control. Feeling abandoned by God and men, they are all of them grotesques of some sort, their freakishness but the existential sign of their alienation. Unable to change their condition, terrified most of the time by what they cannot comprehend, they lash out against it in periodic bouts of violence that are as many rituals of exorcism. After which, broken but somewhat restored, they embark upon another round of mishaps and frustrations that eventually ends up in the same way.

One may not be too far from the mark in stating that all of Crews’s novels so far have been rituals of exorcism, enacting one aspect or another of what René Girard has called “generative violence.” This is a fact of life that the more civilized world does not willingly acknowledge; it was bedrock reality, nevertheless, albeit unperceived as such, in Crews’s world. To Crews, however, between the age of five and six, it asserted itself as an inescapable fact of life, and very consciously so.

I had already learned -without knowing I’d learned it- that every single thing in the world was full of mystery and awesome power. And it was only by right ways of doing things -ritual ways- that kept any of us safe. Making stories about them was not so that we could understand them but so that we could live with them. [...] [F]antasy meant survival” (the end of chapter 6, 90).

Crews’s very act of novel writing is a ritual of exorcism; and each of his novels re-enact a form of it. Perhaps this is why they give such an impression of sameness, despite the abundance and variety of characters and situations. To open a Crews novel is to expect it, to expect freakish characters of some sort involved in violent action of some kind. There are bodybuilders, beauty queens, and karatekas preparing single-mindedly for their competitions; there is a young man eating his car, pound after pound a day; an oversexed Gospel singer given to compulsive fornication; a twentieth century nostalgic of the mediaeval rites of falconry bent on taming a hawk; there is a town of rednecks getting ready to celebrate its annual roundup of rattle snakes. Then there are the freaks, the authentic freaks of the various freak shows, the giants, midgets, fat men and fat ladies, whose deformities are emblematic of the alienated state of mankind.

Unlike as they may appear, they still share the same outlandish extravagance, the same compulsive hankering after something they can’t even name - acceptance, love, happiness , God ... whatever. All of Crews’s characters appear obsessed by something, which may be carnal love, physical perfection, ambition, salvation and penance; and this obsession inevitably involves them in bouts of violence that are as many rituals of exorcism.

Both themes are intricately linked in Crews’s first published novel, The Gospel Singer, in 1968. Condensed, film scenario fashion (in the manner of the scripts Crews himself wrote during his spell with Hollywood), its story might read: Gospel singer’s obsession with female flesh combined with manager’s obsession with penance leads to violent undoing by frenzied mob.

Nor is the spareness of this formulation of the novel’s contents quite alien to the nature of the story. There is no importunate dalliance with irrelevant episodes, no unwarranted sensationalism (as may appear, it must be said, in some of Crews’s lesser novels), no flamboyance for flamboyance sake. The action unfolds in one place and in one day and not so much drifts as it is recklessly and relentlessly driven to its cathartic dénouement. Of all Crews’s novels, The Gospel Singer may be regarded as possessing in the most stringent way the very lineaments of Tragedy: the tension and ineluctability, the hero’s tragic weakness and its implacable retribution, the strict husbanding of means toward end. It also has the crispness of Crews’s best prose, which is not to be found in every one of his novels, but is the prose of A Childhood and of the sequel he is currently writing (judging from the opening chapters that appeared in the Fall 1998 issue of The Southern Quarterly and were reprinted since then in a book of Crews’s interviews, Getting Naked with Harry Crews.)

A surface reading of The Gospel Singer makes it appear for what Crews intended it to be, that is, a Christian tragedy set in our Western world deprived of the old certainties of Christianity, where Gospel singing and religion have turned into a lucrative business. This is heralded by the epigraph Crews added to it: “Men to whom God is dead worship one another,” and further implemented by his specific use of freaks as emblems of man’s estrangement from God. In this respect, Crews shares his fellow Georgia writer Flannery O’Connor’s belief in the essential depravity of men, which both of them make perceptible to their readers through the very visible bodily distortion of the grotesques that people their novels. “If there are freaks in my novels,” Crews admitted in his 1972 interview, “it is only that these people have conditions which are more apparent and more immediate than the people around them. But I am convinced that you and I, all of us, are caught in the same kind of inexplicable, almost blind terror, except that ours is not so apparent” (Getting Naked 30-31). Both Southern writers, however different their faith, have drawn their characters from the same deprived rural neighborhood and, by distorting them into freaks, have made them into “figure[s] of our essential displacement” (Mystery 45).

O’Connor, however, held a staunch belief in the workings, devious as their ways may appear, of God’s Grace, whereas Crews, already as a small boy, has had serious doubts about it. There certainly is no Redeeming God in The Gospel Singer, only bleak humanity engaged in rites and rituals that hark back to a much darker age, when the orgiastic killing of vegetation gods had not yet given rise to its apotropaïc representation on the tragic stage. To the point here, undoubtedly, must be the setting of the novel, the small town named Enigma, and the presence of the eunuchlike, self-flagellating Corybant figure of the Gospel singer’s manager-priest Didymus, among many other references.

Didymus, the nickname of the apostle Thomas, as the Bible-ridden Southern audience may well have known, is also the name of one of the earlier oracles of the god Apollo in Asia Minor before he won his fight against Python and moved to Delphi. Had indeed the Gospel singer kept clear of Enigma and its man-eating Sphinx, Apollo to the end he would have remained, sunny, golden-haired and fair, his Gospel singing progression around the country no less triumphant than Apollo’s progresss around Greece amid his retinue of Muses.

No such glory, however, was to be the novel’s protagonist’s. Enigma erases the radiant image of Apollo as the Gospel singer stumbles into the dark revelries of the god who shared Apollo’s shrine at Delphi, Dionysus. The old Dionysian rites, indeed, loom large over the destructive frenzy of the Enigma mob, while the Gospel singer eventually turns into the sacrificed victim of yet another avatar of the god, i.e. the Dionysus-Zagreus of the Orphic myths.

If the Gospel singer may eventually appear, in the enveloping Christian symbolism of the novel, as a Christ figure in his sacrificial death, he certainly is no innocent victim intended to put an end to the old scapegoating rituals of mankind, as René Girard interprets Christ’s “non sacrifice” on the Cross in his “non sacrificial” reading of the Gospels (in Things hidden since the Foundation of the World). No innocent Savior, the Gospel singer is put to death because of the black sheep he is, lustful and sinful to all excess, and also because the people of Enigma refuse to acknowledge the blackness of their own hearts and transfer it upon him.

All the cogs and wheels of the old scapegoating machinery are set in motion by the end of the novel. Once an adulated member of the community, the Gospel singer is turned into the “inside marginal” (another Girard term) because of the excess of his wealth, fame, and good looks. Magically endowed with exceptional powers by the adoring crowd, the power to heal the lame and the halt, to forestall death, even to bring rain to the parched land, he is made a god, all too human alas! and then murdered for being human and the worst of sinners to boot. In accordance with the motives laid bare in Girard’s book The Scapegoat, he is made to pay for the worst abomination, here his transgression of the strictest taboos of Southern society: Pure Southern Womanhood, with its corollary of racial purity.

Not only does he confess to having perverted the Southern girl who was to be his bride and turned her into an unashamed whore, but he also tells the truth about her killing at the hands of the black preacher who was his childhood pal: not because he has raped her and murdered her to hide the deed, but because, in the words of the Gospel singer, “he wouldn’t screw” her, when she begged for it (Gospel 193).

A seducer, a sinner, and above all a scandalmonger and liar about the sacred tenets of the community, he’s strung with the preacher from the same oak tree by the infuriated crowd, made to atone for Enigma’s own acknowledged suspicion of the black man, its unallowable lust for the beautiful MaryBell, its envy for the preacher who, it is thought, had “had” her, its “mimetic desire” (to use Girard’s seminal expression) for the all-too-successful Gospel singer, and, last of all, for its self-loathing, its despair at its failures and poverty.

Catharsis has been achieved. Law and order are restored to Enigma, but as Roberto Calasso writes in The Ruin of Kasch, which is a reflection on order and disorder, all sacrificial murder is a self-murder, since violence has been allayed for a limited time only, and any member of the same community is liable to become a victim for the next outburst.1

The ritual killing of the pharmacos, in restoring order, also breaches the dis-ruption within the community, re-unites what has been put asunder. In the deadly love-hate embrace which closes The Gospel Singer, Calasso would also recognize the coniunctio which, on the strength of his reading of Vedic texts, he assumes to be the essence of the act of sacrfice, “the ritual foundation of the interweaving of Eros and Thanatos” (Ruin 151).

“I fear my world view is a terribly black, awful one,” Crews admitted in the 1972 interview. “I have only one thing to say to people. As soon as something pleasant and cheerful and confectionery occurs to me, I’ll write about it; but I can only write about whatever comes. And what has come so far has been a kind of blackness” (Getting Naked 31).

Something “confectionery” must undoubtedly have come Crews’s way in the thirty-odd years of his novelwriting: money, recognition, fame, gratification from a work well done. Yet all his novels, to the latest, evince the same bleakness disguised under the trappings of violent farce and baroque action. Reading A Chilhood, however, one can understand why. At five, Crews’s outlook on life was already deeply stamped upon the man he was to become, and rooted into his personality in an ineradicable way. And inescapable too has appeared in his adult life the repetition of the selfsame patterns that have darkened his childhood, as if bred into the bone by family and environment: the violence, outrageous behavior, bodily accidents, hard drinking, losses.

1 Calasso writes: “in offering the sacrifice, we accept–even behind the stratagem of substitution, which temporarily keeps us alive–the fact that we ourselves will one day be devoured, if not by men, then by those gods who are invisible guests at the banquet:

’killing always means killing oneself’” (Ruin 160).

Harry Crews, at sixty-five, seems to have found some peace. Retired from teaching, and from the bottle that has caused so much havoc among writers,^{2} he is currently writing a sequel to A Childhood, starting at the time of his mother’s death.

In one of his latest interviews, Crews has expressed his wonder at being published in France (Getting Naked 340 & 349).3 To what measure of response and success, however, lacking Gallimard’s record of sales, is hard to assess. A first venture in 1974 with the translation of Car (Superbagnole) by Albin Michel had come to nothing. It seems Gallimard has bought the translation rights in the 1990s, reissuing Car under its American title, and a few others of Crews’s novels: La Foire aux Serpents (A Feast of Snakes), La Malédiction du Gitan (The Gypsy’s Curse), Body (same title), Le Chanteur de Gospel (The Gospel Singer), and fortunately, albeit belatedly (in 1997), Des Mules et des Hommes, the French felicitous title for A Childhood, possibly borrowed from, or inspired by, Zora Neale Hurston’s 1935 collection of tales, Mules and Men.

It so happened that the French version of Crews’s “biography of a place” appeared about the same time as the translation of Frank McCourt’s own memoir of his childhood in Ireland, Angela’s Ashes. For those French readers who have read both books, the debate remains open as to whose childhood was the more appalling, whose environment the more unredeemed. In both cases, whether McCourt’s Roman Catholic Ireland or Crews’s Baptist South, combined underpriviledge and ignorance was responsible for their experiencing early in life what a French writer has so adequately called “la maudissure de la vie, ” the everyday sorriness and wretchedness that is like a blight on one’s life. McCourt, however, has not produced any novels so far.

2 cf. James Dickey’s son Christopher’s memoir on his father, Summer of Deliverance.

New York: Simon & Schuster, 1998.

3 “In my life, I finally found an audience. France. The French love me” (340”)… Telling the interviewer that he wished he had The Mulching of America back in order to burn it, he says that “the French just bought it, God bless them… ” adding: “The French are curious people” (349).

The curse, apparently, has never lifted from Crews’s writing life. He, however, has exorcized it in the act of writing, in creating the flamboyant situations and characters of his novels, their farce and drama, and in the better cases, as in The Gospel Singer and A Childhood, the chastening pity and terror of Tragedy.

Bibliography

Erik Bledsoe. “An Interview with Harry Crews,” Southern Quarterly, 37 (Fall 1998); 97-117; reprinted in Getting Naked with Harry Crews, ed. Erik Bledsoe. University Press of Florida, 1999; 331-62.

Roberto Calasso. The Ruin of Kasch, trans. William Weaver & Stephen Sartarelli. Cambridge: Harvard University Press, 1994.

Harry Crews. The Gospel Singer. New York: William Morrow, 1968; the quotations are from the 1969 Dell paperback edition. (Trad. Le chanteur de Gospel. Paris: Gallimard, 1995).

———– . A Childhood. The Biography of a Place. New York: Harper & Row, 1978. (Trad. Des mules et des hommes. Paris: Gallimard, 1997).

———– . “Assault of Memory,” Southern Quarterly, 37 (Fall 1998); 124-138.

Anne Foata. “Interview with Harry Crews,” RANAM, 5, 1972; 207-25; reprinted in Getting Naked with Harry Crews, ed. Erik Bledsoe. Gainesville: University of Florida Press, 1999; 26-48.

René Girard. Violence and the Sacred, trans. Patrick Gregory. Baltimore: Johns Hopkins University Press, 1977.

———— . The Scapegoat, trans. Yvonne Freccero. Baltimore: Johns Hopkins University Press, 1986.

———— . Things Hidden since the Foundation of the World, trans. Stephen Bann & Michael Metteer. Stanford University Press, 1987.

Flannery O’Connor. “Some Aspects of the Grotesque in Southern Fiction,” in Mystery and Manners, ed. Sally and Robert Fitzgerald. New York: Farrar, Straus & Giroux, 1969.

A Childhood, The Gospel Singer, and Writing: Harry Crews’s Rituals of Exorcism

Anne FOATA

L’écriture, pour Harry Crews, cet auteur sudiste né en 1935, est un rite d’exorcisme pour échapper à l’implacable violence et pauvreté qui a marqué son enfance dans un comté du sud de la Géorgie pendant la Dépression. Son autobiographie, A Childhood, qui date de 1978, mais n’a paru en France qu’en 1997 (Des Mules et des Hommes), raconte pourquoi et comment, dès l’âge de cinq ans, il s’est mis à “fabriquer” des histoires à partir du catalogue de ventes par correspondance Sears & Roebuck, une habitude qui devait mener à son premier roman, The Gospel Singer, en